A wooden stick of length 3L is rotated about an end with constant angular velocity w (omega) in a uniform magnetict field B of perpendicular to the plane of motion. If the upper one-third of its length is coated with copper, the potential difference across the whole length of sticks?

Grade:Upto college level

1 Answers

vikas askiitian expert
13 years ago

potential difference will be due to motion of one third part of length...

change in area in 1 revolution, ds = pi(3L)2-pi(2L)2 =5piL2       

angle bw magnetic field & area vector is 0 ..

flux passing through the one third part of length over 1 revolution is

@ = Bdscos0        (flux)

@ = Bds

e = d@/dt = Bds/dt =B5piL2/T

T=2pi/W

e = B5L2W/2